OpenStudy (anonymous):

Simplify 3 square root of 128 please (:

OpenStudy (tkhunny):

Ever hear of a "Prime Factorization"?

OpenStudy (wolf1728):

128 = 2^7 128 = 2*64 square root of 128 = 8 * square root of 2 So 3 * square root of 2 = 24 * square root of 2